# hyperrefToken not allowed in a PDF string

### v1

@@ Appendix A Proofs - Not polished
Proof  As we recalled in the proof of Lemma 1, Chowdhury and Gopalan (2017) have shown that with a choice of $\beta_{n}=B+R\sqrt{2\mathopen{}\mathclose{{}\left(\ln(e/\delta)+\gamma_{n}}\right)}$, it holds with probability of at least $1-\delta$ that $s(\bm{x})\in\mathopen{}\mathclose{{}\left[\mu_{n}^{s}(\bm{x})\pm\beta_{n}\sigma_{n}^{s}(\bm{x})}\right]$ for all $\bm{x}\in\mathcal{X}$ and for all $n$. If we now take a GP prior and a dataset $\mathcal{D}$ such that $\mu_{n}$ and $\sigma_{n}$ define the posterior $GP$ obtained by conditioning on the subset of $\mathcal{D}$ $\mathcal{D}_{n}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, then the result by Chowdhury and Gopalan (2017) means that, with probability of at least $1-\delta$, $s(\bm{x})\in\mathopen{}\mathclose{{}\left[\mu_{n}^{s}(\bm{x})\pm\beta_{n}\sigma_{n}^{s}(\bm{x})}\right]$ for all $\bm{x}$ and for all $n\leq\mathinner{\!\mathopen{}\mathclose{{}\left\lvert\mathcal{D}}\right\rvert}$.
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}\mathrel{\mathop{:}}g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (23)
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}\mathrel{\mathop{:}}g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\}$ (27)
@@
[…]
@@
[…]
@@
Lastly, before we can prove Theorems 1 and 2, we need two more results about $\bar{S}(s)$, which we recall we defined as the total true safe set: $\bar{S}(s)\coloneqq\mathopen{}\mathclose{{}\left\{\bm{x}\in\mathcal{X}\mathrel{\mathop{:}}s(\bm{x})\geq 0}\right\}$.
@@
[…]
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}\mathrel{\mathop{:}}\beta_{n}g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\}$ (29)

### v2

@@ Appendix A Proofs - Not polished
[…]
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}:g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (23)
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}:g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\}$ (27)
@@
$R^{\text{GP}_{\mathcal{D}}}_{\varepsilon}(\Omega)=\{\bm{x}\in\mathcal{X}:\bm{x}\in S_{\text{GP}_{\mathcal{D}}}\}.$ (28)
@@
Let $\Omega$ be a subset of the domain, then for any GPD, GPD $\in$ $\bm{\beta}\text{-GP}_{f}^{\varepsilon}(\Omega)$ implies the inclusion $R^{\text{GP}_{\mathcal{D}}}_{\varepsilon}(\Omega)\supseteq R_{\varepsilon}(\Omega)$.
Proof  The claim follows directly by the definitions of the two operators $R^{\text{GP}_{\mathcal{D}}}_{\varepsilon}$ and $R_{\varepsilon}$, since a parameter $\bm{x}$ that is not classified as safe by the specific GPD is not, by definition, part of $R_{\varepsilon}(\Omega)$.
@@
Lastly, before we can prove Theorems 1 and 2, we need two more results about $\bar{S}(s)$, which we recall we defined as the total true safe set: $\bar{S}(s)\coloneqq\mathopen{}\mathclose{{}\left\{\bm{x}\in\mathcal{X}:s(\bm{x})\geq 0}\right\}$.
@@
For any GPD such that $h(\bm{x})\in[\mu_{n}(\bm{x})\pm\beta_{n}\sigma_{n}(\bm{x})]$ for all $\bm{x}$ and for all $n\leq|\mathcal{D}|$, it holds that $S_{n}\subseteq\bar{S}(s)$.
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}:\beta_{n}g^{-1}\mathopen{}\mathclose{{}\left(\frac{C_{1}\gamma_{N}}{N}}\right)\leq\varepsilon}\right\}$ (29)